OpenStudy (anonymous):

The current in an electric circuit of constant voltage varies inversely as the resistance. When the current is 30 amps, the resistance is 20 ohms. Find the current when the resistance is 25 ohms.

OpenStudy (anonymous):

24 amp

OpenStudy (anonymous):

V=IR, voltage equals current times resistance. So, V=30*20=600 volts, then. your new current I=V/R=600/25=24 amp.

OpenStudy (anonymous):

in an inverse relation xy = k in first case 30 * 20 = 600 in second case , current * 25 = 600 => current = 600/25 = 24 amp